OpenStudy (anonymous):

If you graph f(x)= x^2 + y^2 then it is a circle but then how would you graph the square root of f(x)?

OpenStudy (anonymous):

how would you go about graphing it*

OpenStudy (anonymous):

but f(x)= x^2 + y^2 isn't a circle

OpenStudy (anonymous):

i'm pretty sure f(x)= x^2 + y^2 is a circle... the square root of f(x) is half a circle if it's positive sqrt of f(x), it'll be the top half of the circle if it's negative, it will be the bottom half.

OpenStudy (anonymous):

what would be the radius?

OpenStudy (anonymous):

good point, i didn't catch that uzma right but i think the point of the question is to understand that the square roots is half a circle so you need a value for f(x), otherwise, as uzma said, it's not a circle

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

it must be x^2 + y^2= costant, then centre is (0,0) and radius is sqrt(constant)

OpenStudy (anonymous):

or we can have the general eq x^2 +y^2 +2fx+2gy +c=0